Find the center and radius of the circle:
x^2+y^2-6x+2y+6=0

Answer by stanbon(57984)

Find the center and radius of the circle:
x^2+y^2-6x+2y+6=0
-----
Complete the square on the x-terms and on the y-terms:
----
x^2-6x+9 + y^2+2y+1 = -6+9+1
----
(x-3)^2 + (y+1)^2 = 4
----
Center: (3,-1)
Radius: sqrt(4) = 2
=====================
Cheers,
Stan H.